OpenStudy (anonymous):

Nina can ride her bike 63,360 feet in 3,400 seconds, and Sophia can ride her bike 10 miles in 1 hour. What is Nina's rate in miles per hour if there are 5,280 feet in a mile? Which girl bikes faster?

OpenStudy (tkhunny):

Dimensional Analysis \(\dfrac{63360\;ft}{3400\;sec}\cdot\dfrac{1\;mile}{5280\;ft}\cdot\dfrac{3600\;sec}{1\;hr} = \) What \(\dfrac{miles}{hour}\)?

OpenStudy (anonymous):

Thanks!

OpenStudy (tkhunny):

What did you get? Can you do other transformations?

OpenStudy (anonymous):

\[\frac{63360}{\frac{5280\ 3400}{3600}}=\frac{216}{17}=12.7059 \text{ mph} \] Sorry, LaTex is not reproducing the text that it was fed.

OpenStudy (anonymous):

I got 12.7. And it was one of my answer choices,

OpenStudy (anonymous):

Thank you too robtobey!

OpenStudy (anonymous):

Very good. Congratulations.